Find the line y=mx through the origin, with positive slope, which together with the parabola y=x^2 encloses a region of area 4/3.

1
Expert's answer
2021-11-16T14:28:12-0500

Find the point(s) of intersection


"mx=x^2"

"x_1=0, x_2=m"

"Point(0,0)" and "Point(m, m^2)"


If "m>0"


"Area=A=\displaystyle\int_{0}^{m}(mx-x^2)=[\dfrac{mx^2}{2}-\dfrac{x^3}{3}]\begin{matrix}
 m \\
 0
\end{matrix}"

"=\dfrac{m^3}{2}-\dfrac{m^3}{3}=\dfrac{m^3}{6} ({units}^2)"

Substitute


"\dfrac{m^3}{6}=\dfrac{4}{3}"

"m=2"